OpenStudy (anonymous):

Which equation represents the line whose slope is -2 and that passes through point (0, 3)?

OpenStudy (anonymous):

y-y1=m(x-x1) y-3= -2(x-0) y-3=-2x 2x+y=3

OpenStudy (amonoconnor):

Plug it into Slope-Intercept Form (y = mx + b), solving for "b". Then, plug in m with that found value to determine the equation. So... (3) = (-2)(0) + ... ?